OpenStudy (anonymous):

What is the frequency of electromagnetic energy with a wavelength of 5 x 10^-7 meters?

OpenStudy (michele_laino):

we have to apply this formula: \[\Large \nu = \frac{c}{\lambda } = \frac{{3 \times {{10}^8}}}{{5 \times {{10}^{ - 7}}}}\]

OpenStudy (michele_laino):

where \nu is the requested frequency

OpenStudy (anonymous):

ok! so we get 6E14 ?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

is that our final solution? :O

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ooh okie yay! thank you!!:D

OpenStudy (michele_laino):

:)